OpenStudy (anonymous):

Write in scientific notation 5,800,000,000,000/0.000029

OpenStudy (rulnick):

2x10^17

OpenStudy (anonymous):

\[5.8\times 10^{12}\div2.9\times 10^{-5}\] divide the how part, subtract the exponents to get \[2\times 10^{12-(-5)}=2\times 10^{17}\]